# On the hydrogen atom in an spherical box

### v1

@@ 2 The model
where $m_{e}$ is the electron mass and $K>0$ is the strength of the Coulomb potential (with units of energy×$\times$×length). For simplicity, we assume that the nucleus is clamped at the origin. The solutions $\psi(r,\theta,\phi)$ in spherical coordinates $x=r\sin\theta\cos\phi$, $y=r\sin\theta\sin\phi$, $z=r\cos\theta$, $0<\theta<\pi$, $0\leq\phi<2\pi$, satisfy the boundary condition $\psi\left(r_{0},\theta,\phi\right)=0$ because of the impenetrable wall of a spherical box of radius $r_{0}$. Therefore, $0<r\leq r_{0}$.
In order to discuss some useful analytical properties of the solutions to the eigenvalue equation it is convenient to carry out the scaling transformation $(x,y,z)\rightarrow(L\tilde{x},L\tilde{y},L\tilde{z})$, $r\rightarrow L\tilde{r}$, $\nabla^{2}\rightarrow L^{-2}\tilde{\nabla}^{2}$, where $L$ is a real positive constant with units of length. In this way, we derive a useful dimensionless eigenvalue equation $\tilde{H}\tilde{\psi}=\tilde{E}\tilde{\psi}$, where[8]
$\tilde{H}=\frac{m_{e}L^{2}}{\hbar^{2}}H=-\frac{1}{2}\tilde{\nabla}^{2}-\frac{m_{e}LK}{\hbar^{2}\tilde{r}}.$ (2)
The boundary condition becomes $\tilde{\psi}\left(\tilde{r}_{0},\theta,\phi\right)=0$, where $\tilde{r}_{0}=r_{0}/L$, and the unit of energy is $\hbar^{2}/\left(m_{e}L^{2}\right)$. If $E\left(r_{0},K\right)$ denotes an eigenvalue, then equation (2) tells us that
@@
The scaling transformation of the Hamiltonian operator proves useful for the derivation of many properties of quantum-mechanical systems as discussed elsewhere[8]. In what follows we add another example.
If we choose $L=r_{0}$ then
$\tilde{H}=-\frac{1}{2}\tilde{\nabla}^{2}-\frac{m_{e}r_{0}K}{\hbar^{2}\tilde{r}},$ (4)
and $\tilde{\psi}(1,\theta,\phi)=0$. Therefore,
$\lim\limits_{r_{0}\rightarrow 0}\tilde{H}=-\frac{1}{2}\tilde{\nabla}^{2},$ (5)
and the problem reduces to a particle within a spherical box of radius $\tilde{r}_{0}=1$. In this case the unit of energy is $\hbar^{2}/\left(m_{e}r_{0}^{2}\right)$.
Alternatively, we may also choose $L=\hbar^{2}/(m_{e}K)$ that leads to
$\tilde{H}=-\frac{1}{2}\tilde{\nabla}^{2}-\frac{1}{\tilde{r}},$ (6)
and the unit of energy $\hbar^{2}/(m_{e}L^{2})=m_{e}K^{2}/\hbar^{2}$.
Throughout this paper we will omit the tilde over the dimensionless quantities and will write, for example, the Hamiltonian operator (4) as
$H=-\frac{1}{2}\nabla^{2}-\frac{\beta}{r},\;\beta=\frac{m_{e}r_{0}K}{\hbar^{2}}.$ (7)
Before proceeding with our discussion of the model (7) we want to point out that expressions like “we choose units such that $\hbar=m=1$” are meaningless if we do not clearly indicate the units of length and energy actually used. Note that the two dimensionless Hamiltonians $\tilde{H}$ shown in equations (4) and (6) formally correspond to setting $\hbar=m=1$ though the units of length and energy in one case are different from those in the other.
If we substitute the two values of $L$ shown above into equation (3) and equate the results we can easily derive the relationships
$\tilde{E}(1,\beta)$ $=$ $\beta^{2}\tilde{E}(\beta,1),$
$\tilde{E}\left(1,\tilde{r}_{0}\right)$ $=$ $\tilde{r}_{0}^{2}\tilde{E}\left(\tilde{r}_{0},1\right),$ (8)
that clearly show the connection between $\beta$ and $r_{0}$ in the two alternative approaches to the problem.
The Schrödinger equation for any of the Hamiltonian operators shown above is separable in spherical coordinates as $\psi(r,\theta,\phi)=R(r)Y_{l}^{m}$, where $l=0,1,\ldots$ and $m=0,\pm 1,\pm 2,\ldots,\pm l$ are the angular and magnetic quantum numbers, respectively, and $Y_{l}^{m}$ are the well-known spherical harmonics. The energy eigenvalues depend only on the radial quantum number $n=0,1,\ldots$ and on $l$ so that we write them as $E_{nl}$ from now on. For convenience, we do not resort to the principal quantum number $n_{p}=n+l+1=1,2,\ldots$ that is mostly useful in the case of the free hydrogen atom. It follows from the Hamiltonian operator (6) that
$\lim\limits_{r_{0}\rightarrow\infty}E_{nl}=E_{nl}^{H}=-\frac{1}{2(n+l+1)^{2}},$ (9)
from which we conclude that $E_{nl}^{H}>E_{n^{\prime}l^{\prime}}^{H}$ if $n+l>n^{\prime}+l^{\prime}$. This obvious inequality will be useful later on.
Since $E(1,0)$ is positive and $\lim\limits_{\beta\rightarrow\infty}\beta^{-2}E(1,\beta)$ is negative, then for each eigenvalue $E_{nl}$ there is a value $\beta=\beta_{nl}^{c}$ such that $E_{nl}\left(\beta_{nl}^{c}\right)=0$. We will discuss these critical values of $\beta$ in section 4.
@@ 3 Exact polynomial solutions
The radial part of the Schrödinger equation for the Hamiltonian operator (7) is
$-\frac{1}{2}\left(\frac{d^{2}}{dr^{2}}+\frac{2}{r}\frac{d}{dr}\right)R(r)+\left[\frac{l(l+1)}{2r^{2}}-\frac{\beta}{r}\right]R(r)=ER(r),$ (10)
@@
where we have set $E=-\alpha^{2}/2$ in order to remove one of the terms.
In order to obtain exact polynomial solutions we require that $c_{\nu}\neq 0$ and $c_{\nu+1}=c_{\nu+2}=0$, $\nu=0,1,\ldots$. These conditions are satisfied if $B_{\nu}=0$ from which we obtain
@@
The expression for $E$ in equation (13) does not give us the spectrum of the problem. Note that $E=0$ when $\beta=0$ while the Hamiltonian (7) tells us that we should obtain the spectrum of the particle in a box of radius $r_{0}=1$ when $\beta=0$. Besides, the polynomial solutions only provide negative eigenvalues while all the eigenvalues are positive for sufficiently small values of $\beta$ as argued in section 2. Any smart reader may think that it is not necessary to stress such an obvious fact but unfortunately many researchers have misinterpreted the polynomial solutions of several conditionally-solvable models as discussed elsewhere[11, 12].
Since $B_{\nu}=0$ the only remaining condition is $c_{\nu+1}=0$ from which we obtain $\nu+1$ roots $\beta_{l}^{(\nu,i)}$, $i=0,1,,\ldots,\nu$, that we arbitrarily arrange so that $\beta_{l}^{(\nu,i+1)}>\beta_{l}^{(\nu,i)}$. Thus, the energies of the polynomial solutions should be more properly written as
@@
In the expressions above $\nu$ is the degree of the polynomial factor in the exact solution and $i$ is the number of real zeros of the polynomial in the interval $10<r<$ 1. For this reason $i$ (and not $\nu$) is related to the radial quantum number $n$. This fact was overlooked by many researchers as discussed in the papers just mentioned[11, 12]. Of particular interest are the roots
$\beta_{l}=\beta_{l}^{(0,1)}=(l+1)(l+2),\;E_{1}^{(0,1)}=-\frac{(l+1)^{2}}{2},$ (16)
@@ 4 Accurate numerical results
One can obtain accurate numerical results in several ways as shown in suitable reviews on the subject[1, 2, 3, 4]. Here, we resort to the RRM[6, 7] that provides increasingly accurate upper bounds to the exact eigenvalues[13, 14].
For simplicity, we resort to the non-orthogonal basis set
@@
The RRM secular equations are well-known[6, 7, 14, 15] and will not be discussed here.
Table 1 shows some eigenvalues for the case $\beta=0$. We appreciate that there are several cases in which $E_{nl}^{PB}>E_{n^{\prime}l^{\prime}}^{PB}$ with $n+l<n^{\prime}+l^{\prime}$. Consequently, we expect that such eigenvalues $E_{nl}$ and $E_{n^{\prime}l^{\prime}}$ should cross at some nonzero value of $\beta$ because $E_{n^{\prime}l^{\prime}}^{H}>E_{nl}^{H}$ as argued in section 2.
Figure 1 shows the lowest eigenvalues with $l=0,1,2,3$. We appreciate the crossings at $\beta=\beta_{0}=2$ between $E_{10}$ and $E_{02}$ and also between $E_{20}$ and $E_{12}$. This fact suggests that the values $\beta_{l}$ of the model parameter given by the exact polynomial solutions are special. It is worth noting that the former accidental degeneracy appeared in an earlier paper[16] (see also table 4 in page 140 in reference[2]) but nobody paid attention to it as far as we know. The blue points in figure 1 are values of exact energies given by equation (15) when $i=0$. Since the polynomial factors of such solutions do not exhibit nodes, then they correspond to the ground state as the figure already shows.
@@
At present we are unable to prove this conjecture but all our numerical results confirm it.
The RRM enables us to obtain the critical values of $\beta$ introduced in section 2. We simply set $E=0$ in the secular equation[6, 7, 14, 15] and solve for $\beta$. Table 3 shows some critical values of $\beta$ for $l=0,1,2,3$. The roots $\beta_{l}^{(\nu,i)}$ decrease as $\nu$ increases and appear to approach to $\beta_{nl}^{c}$ when $n=i$. This fact, which is shown in figure 2, suggests the following
Conjecture 2
$\lim\limits_{\nu\rightarrow\infty}\beta_{l}^{(\nu,n)}=\beta_{nl}^{c}$
At present we are unable to prove this conjecture that our numerical results confirm.
@@ 5 Conclusions
In this paper we have shown several aspects of the well known hydrogen atom inside a box with an impenetrable spherical wall that have passed unnoticed, as far as we know. In the first place, a suitable scaling of the Hamiltonian operator is extremely useful for the derivation of several general properties of the eigenvalues. In the second place, the radial part of the Schrödinger equation is conditionally solvable. In the third place, there are most interesting seemingly accidental degeneracies that take place at particular values of the box radius, some of which can be determined from the conditionally-solvable condition. In the fourth place, some of the roots stemming from the conditionally-solvable condition appear to converge towards the critical values of the model parameter. At present we cannot prove the two latter results rigorously and have, therefore, presented them as conjectures. In this analysis the RRM proved to be most useful.

### v3

# Exact solutions, critical parameters and accidental degeneracy for the hydrogen atom in a spherical box

@@ 2 The model
[…]
The dimensionless box radius is $\tilde{r}_{0}=r_{0}/L$. If $E\left(r_{0},K\right)$ denotes an eigenvalue of $H$, then this equation tells us that
@@
If $L=r_{0}$ we have
$E\left(r_{0},K\right)=\frac{\hbar^{2}}{m_{e}r_{0}{}^{2}}E\left(1,\beta\right),\;\beta=\frac{m_{e}r_{0}K}{\hbar^{2}}.$ (4)
On the other hand, if $L=\hbar^{2}/\left(m_{e}K\right)$ we have
$E\left(r_{0},K\right)=\frac{\hbar^{2}}{m_{e}r_{0}{}^{2}}\beta^{2}E\left(\beta,1\right).$ (5)
It follows from equations (4) and (5) that
$E(1,\beta)=\beta^{2}E(\beta,1).$ (6)
It is clear that $E(1,\beta)$ is the dimensionless energy of a hydrogen atom with interaction $\beta/r$ and unit box radius, while $E(\beta,1)$ is the dimensionless energy of a hydrogen atom with interaction 1/r-1/r and a dimensionless box radius equal to $\beta$. Both descriptions of the problem are related by the simple expression (6).
The Schrödinger equation in any of the two cases discussed above is separable in spherical coordinates as $\psi_{nlm}(r,\theta,\phi)=R_{nl}(r)Y_{l}^{m}(\theta,\phi)$, where $n=0,1,\ldots$, $l=0,1,\ldots$ and $m=0,\pm 1,\pm 2,\ldots,\pm l$ are the radial, angular and magnetic quantum numbers, respectively, and $Y_{l}^{m}$ are the well-known spherical harmonics. The energy eigenvalues depend only on $n$ and $l$ so that we write them as $E_{nl}$ from now on. For convenience, we do not resort to the principal quantum number $n_{p}=n+l+1=1,2,\ldots$ that is mostly useful in the case of the free hydrogen atom.
It is clear that $E_{nl}(1,0)$ are the eigenvalues of a free electron in a box of unit radius; therefore, they are all positive. On the other hand,
$\lim\limits_{\beta\rightarrow\infty}E_{nl}(\beta,1)=\lim\limits_{\beta\rightarrow\infty}\beta^{-2}E_{nl}(1,\beta)=E_{nl}^{H}=-\frac{1}{2(n+l+1)^{2}},$ (7)
are the dimensionless energies of the free atom. In this case $E_{nl}^{H}>E_{n^{\prime}l^{\prime}}^{H}$ if $n+l>n^{\prime}+l^{\prime}$. This obvious inequality will be useful later on.
Since $E_{nl}(1,0)=E_{nl}^{PB}>0$ and $\lim\limits_{\beta\rightarrow\infty}\beta^{-2}E_{nl}(1,\beta)<0$, then for each eigenvalue $E_{nl}$ there is a value $\beta=\beta_{nl}^{c}$ such that $E_{nl}\left(\beta_{nl}^{c},1\right)=E_{nl}\left(1,\beta_{nl}^{c}\right)=0$. We will calculate some of these critical values of $\beta$ in section 4.
@@ 3 Exact polynomial solutions
The radial part of the Schrödinger equation for the dimensionless Hamiltonian operator
$H=-\frac{1}{2}\nabla^{2}-\frac{\beta}{r},$ (8)
is
$\mathcal{H}R(r)=ER(r),\;\mathcal{H}=-\frac{1}{2r^{2}}\frac{d}{dr}r^{2}\frac{d}{dr}+\frac{l(l+1)}{2r^{2}}-\frac{\beta}{r},$ (9)
@@
if $E=-\alpha^{2}/2$.
In order to obtain exact polynomial solutions we require that $c_{\nu}\neq 0$ and $c_{\nu+1}=c_{\nu+2}=0$, $\nu=0,1,\ldots$. These conditions lead to $B_{\nu}=0$ from which we obtain
@@
The expression for $E$ in equation (12) does not give us the spectrum of the problem. Note that $E=0$ when $\beta=0$ while the Hamiltonian (8) tells us that we should obtain the spectrum of the particle in a box of radius $r_{0}=1$ when $\beta=0$. Besides, the polynomial solutions only provide negative eigenvalues while all the eigenvalues of the model are positive for sufficiently small values of $\beta$ as argued in section 2. Any smart reader may think that it is not necessary to stress such an obvious fact but unfortunately many researchers have misinterpreted the polynomial solutions of several conditionally-solvable models as discussed elsewhere[11, 12].
Since $B_{\nu}=0$ the only remaining condition is $c_{\nu+1}=0$ from which we obtain $\nu+1$ roots $\beta_{l}^{(\nu,i)}$, $i=0,1,,\ldots,\nu$, that we conveniently arrange so that $\beta_{l}^{(\nu,i+1)}>\beta_{l}^{(\nu,i)}$. Thus, the energies of the polynomial solutions should be more properly written as
@@
In the expressions above $\nu$ is the degree of the polynomial factor of the exact solution (10) and one can verify that $i$ is the number of real zeros in the interval $0<r<1$. For this reason, $i$ (and not $\nu$) is the radial quantum number $n$. This fact was overlooked by many researchers as discussed in the papers just mentioned[11, 12].
Since $i=n$ we conclude that $E_{nl}\left(1,\beta_{l}^{(\nu,n)}\right)=E_{l}^{(\nu,n)}<0$. The Hellmann-Feynman theorem[17, 18]
$\frac{dE_{nl}(1,\beta)}{d\beta}=-\left\langle\frac{1}{r}\right\rangle_{nl},$ (15)
tells us that $E_{nl}$ decreases with $\beta$. Since $E_{nl}\left(1,\beta_{nl}^{c}\right)=0$ we conclude that $\beta_{l}^{(\nu,n)}>\beta_{nl}^{c}$. Numerical results show that $\beta_{l}^{(\nu,n)}$ decreases with $\nu$ as shown in Table 1 for $\beta_{0}^{(\nu,n)}$, $n=0,1,2,3$. From these results and $\lim\limits_{\nu\rightarrow\infty}E_{l}^{(\nu,n)}=0$ we may reasonably put forward the following
Conjecture 1
$\lim\limits_{\nu\rightarrow\infty}\beta_{l}^{(\nu,n)}=\beta_{nl}^{c}$
In section 4 we will show numerical results that support this conjecture. Of particular interest are the roots
$\beta_{l}=\beta_{l}^{(0,0)}=(l+1)(l+2),\;E_{l}^{(0,0)}=-\frac{(l+1)^{2}}{2},$ (16)
@@ 4 Accurate numerical results
One can obtain accurate numerical energies for the hydrogen atom in an spherical box in several ways as shown in suitable reviews on the subject[1, 2, 3, 4]. Here, we resort to the RRM[6, 7] that provides increasingly accurate upper bounds to the exact eigenvalues[13, 14].
For simplicity, we choose the non-orthogonal basis set
@@
The RRM secular equations are well-known[6, 7, 14, 15] and we will just outline them in what follows. In order to solve the radial equation (9) we propose and ansatz of the form
$\varphi(r)=\sum_{i=0}^{N-1}c_{i}f_{il}(r),$ (18)
and the RRM leads to the secular equation
$\mathbf{Hc}=W\mathbf{Sc},$ (19)
where $\mathbf{H}$ and $\mathbf{S}$ are $N\times N$ matrices with elements $H_{ij}=\left\langle f_{il}\right|\mathcal{H}\left|f_{jl}\right\rangle$ and $S_{ij}=\left\langle f_{il}\right.\left|f_{jl}\right\rangle$, respectively, and $\mathbf{c}$ is a $N\times 1$ column vector with elements $c_{i}$. The approximate eigenvalues $W_{nl}$, $n=0,1,\ldots,N-1$, are roots of the secular determinant $\left|\mathbf{H}-W\mathbf{S}\right|=0$. They approach the exact eigenvalues $E_{nl}$ from above which facilitates the estimation of the accuracy of the calculation. In the present case
$\left\langle f\right.\left|g\right\rangle=\int_{0}^{1}f(r)g(r)r^{2}\,dr.$ (20)
Table 2 shows some eigenvalues $E_{nl}^{PB}=E_{nl}(1,0)$. We appreciate that there are several cases in which $E_{nl}^{PB}<E_{n^{\prime}l^{\prime}}^{PB}$ when $n+l>n^{\prime}+l^{\prime}$. Consequently, we expect that such eigenvalues $E_{nl}$ and $E_{n^{\prime}l^{\prime}}$ should cross at some nonzero value of $\beta$ because $E_{nl}^{H}>$ $E_{n^{\prime}l^{\prime}}^{H}$ as argued in section 2.
Figure 1 shows the lowest eigenvalues with $l=0,1,2,3$. We appreciate the crossings at $\beta=\beta_{0}=2$ between $E_{10}$ and $E_{02}$ and also between $E_{20}$ and $E_{12}$. This fact suggests that the values $\beta_{l}$ (16) of the model parameter given by the truncation condition are special. It is worth noting that the former accidental degeneracy at $\beta=2$ appeared in an earlier paper[16] (see also table 4 in page 140 in reference[2]) but nobody paid attention to it as far as we know. The blue points in figure 1 are values of exact energies given by equation (14) when $i=0$. Since the polynomial factors of such solutions do not exhibit nodes, then they correspond to the ground state as the figure already shows.
@@
At present we are unable to prove this conjecture rigorously.
The RRM enables us to obtain the critical values of $\beta$ introduced in section 2. We simply set $E=0$ in the secular equation and solve for $\beta$. Table 4 shows some critical values of $\beta$ for $l=0,1,2,3$. As discussed in section 3 the roots $\beta_{l}^{(\nu,n)}$ approach to $\beta_{nl}^{c}$ from above when $\nu$ increases. Figure 2, illustrates the rate of convergence.
@@ 5 Conclusions
In this paper we have shown several aspects of the well known hydrogen atom inside a box with an impenetrable spherical wall that have passed unnoticed, as far as we know. In the first place, a suitable scaling of the Hamiltonian operator is extremely useful for the derivation of several general properties of the eigenvalues. In the second place, the radial part of the Schrödinger equation is conditionally solvable. In the third place, there are most interesting seemingly accidental degeneracies that take place at particular values of the box radius that are determined by a truncation condition. In the fourth place, some of the roots given by the truncation condition appear to converge towards the critical values of the model parameter. At present we cannot prove the two latter results rigorously and have, therefore, presented them as conjectures. In this analysis the RRM proved to be most useful.